OpenStudy (wolfman77):

Quick help and stuff please!! Will Medal and Fan!! Which statement comparing x² + 10 and x + x + x is true when x=−5 ? x2+10x+x+x x² + 10 = x + x + x

OpenStudy (evoker):

Plug in -5 into each slot, note for the first case the negative will be squared also.

TheSmartOne (thesmartone):

x^2 + 10 Plug in x = -5 (-5)^2 + 10 = ?? x + x + x Plug in x = -5 (-5) + (-5) + (-5) = ??

OpenStudy (wolfman77):

= -15

OpenStudy (wolfman77):

?

TheSmartOne (thesmartone):

and what is (-5)^2 + 10 = ?

OpenStudy (wolfman77):

35?

TheSmartOne (thesmartone):

correct so which is greater? x^2 + 10 or x + x + x

OpenStudy (wolfman77):

the first one

TheSmartOne (thesmartone):

correct and the greater sign looks like this > so your answer is x^2+10>x+x+x

OpenStudy (wolfman77):

Yes. Thanks!!

TheSmartOne (thesmartone):

Anytime! :)